OpenStudy (anonymous):

integral sqrt(((1-cos(x))/2))

OpenStudy (raden):

too many bracket there :p can u figure out of it

OpenStudy (anonymous):

integral sqrt((1-cos(x))/2)

OpenStudy (raden):

hint : sqrt((1-cos(x))/2) = sin(x/2)

OpenStudy (raden):

use the basic formula : int (sinAx) = -cos(Ax)/A

OpenStudy (anonymous):

oh nice thank you!

OpenStudy (raden):

what ur answer, show to me for verify

OpenStudy (anonymous):

-2cos(x/2)+c

OpenStudy (raden):

yes, that's right

OpenStudy (anonymous):

integral 4tan^3(x)dx ?

OpenStudy (raden):

tan^3 x = tan^2 x tanx = (sec^2 x -1)tanx = sec^2 x tanx - tanx then, integrate

OpenStudy (anonymous):

how do you integrate the last part?

OpenStudy (raden):

case I : int(sec^2 x tanx) dx let u=tanx du=sec^2 x dx it can be int(u) du = 1/2 * u^2 = 1/2 * tan^2 x

OpenStudy (raden):

case II ; - int(tanx) dx = -(sinx/cosx) dx let u=cosx du = -sinx dx so, - int(tanx) dx = 1/u du = ln(u) = ln(cosx)

OpenStudy (raden):

so, the answer is 1/2 * tan^2 x + ln(cosx) + c

OpenStudy (raden):

i hope this make sense for u :*

OpenStudy (anonymous):

it does thank you!

OpenStudy (anonymous):

will you be available tomorrow?

OpenStudy (raden):

welcome... not sure if u have problems u can tag me or inbox, so next time i can read it

OpenStudy (raden):

sorry, i want to play table tennis now.... byee

OpenStudy (anonymous):

thanks. have fun!

OpenStudy (anonymous):

integral 25cos^2(x)

OpenStudy (raden):

sorry... i just new online use identity : cos^2 x = (1+cos2x)/2 so, int (25cos^2 x) dx = int (25(1+cos2x)/2) dx = int (25/2 + (25cos2x)/2) dx = (25/2) x + (25/4) sin2x + c because sin2x = 2sinxcosx it also can be = (25/2) x + (25/4) 2sinxcosx + c = (25/2) x + (25/2)sinxcosx + c